# DiaCDM: Cognitive Diagnosis in Teacher-Student Dialogues using the Initiation-Response-Evaluation Framework

### v1

@@ 1 Introduction
CDMs aim to infer students’ mastery of knowledge components, and existing approaches are mainly divided into traditional psychometric methods and machine learning methods. The former, represented by IRT [17] and its multidimensional extensions, characterize students’ cognitive states through latent variables; the latter leverage deep learning to capture complex student–item interactions and further incorporate multimodal information, such as Neural Cognitive Diagnosis Model (NCDM) [8] and Knowledge-Association Neural Cognitive Diagnosis Model (KaNCD) [18], as well as DIRT [19] and NCD+ [8]. However, these methods are mostly limited to static, test-based settings and are difficult to adapt to the dynamic interactions of real teaching. Teacher-student dialogue, as a typical interactive scenario, usually follows the IRE (Initiation–Response–Evaluation) structure [20]. Teachers’ questions and feedback, as well as students’ responses, contain rich and diverse diagnostic signals, and studies have shown that high-quality dialogues can effectively reveal and influence students’ cognitive processes. Nevertheless, these potential diagnostic signals have not been fully utilized in existing dialogue-based tutoring systems, which often focus more on dialogue generation and interaction strategies, lacking systematic and fine-grained student modeling. Therefore, extending CDMs to dialogue settings and achieving cognitive diagnosis from dynamic interactions remain important but challenging open problems.
@@ 5 Experiments
Datasets. We conducted experiments using three datasets: CoMTA, MathDial, and PMTD. These datasets cover human-AI and real teacher-student dialogue interactions, demonstrating the applicability of our approach across diverse dialogue setting.

### v3

@@ 1 Introduction
CDMs aim to infer students’ mastery of knowledge components, and existing approaches are mainly divided into traditional psychometric methods and machine learning methods. The former, represented by IRT [17] and its multidimensional extensions, characterize students’ cognitive states through latent variables; the latter leverage deep learning to capture complex student–item interactions and further incorporate multimodal information, such as Neural Cognitive Diagnosis Model (NCDM) [8] and Knowledge-Association Neural Cognitive Diagnosis Model (KaNCD) [18], as well as DIRT [19] and NCD+ [8]. However, these methods are mostly limited to static, test-based settings and are difficult to adapt to the dynamic interactions of real teaching. Teacher-student dialogue, as a typical interactive scenario, usually follows the IRE (Initiation–Response–Evaluation) structure [20]. Teachers’ questions and feedback, as well as students’ responses, contain rich and diverse diagnostic signals, and studies have shown that high-quality dialogues can effectively reveal and influence students’ cognitive processes. Nevertheless, these potential diagnostic signals have not been fully utilized in existing dialogue-based tutoring systems, which often focus more on dialogue generation and interaction strategies, lacking systematic and fine-grained student modeling. Therefore, extending CDMs to dialogue scenarios and achieving cognitive diagnosis from dynamic interactions remains an important yet challenging research direction. Although recent studies have attempted to incorporate knowledge tracing tasks into teacher–student dialogue settings [21], achieving accurate cognitive diagnosis in open-ended interactions still poses a significant challenge for CDM models.
@@ 5 Experiments
Datasets. Following [21], we conducted experiments using three datasets: CoMTA, MathDial, and PMTD. These datasets cover human-AI and teacher-student dialogue interactions, demonstrating the applicability of our approach across diverse dialogue settings.